OpenStudy (anonymous):

Its two questions. I need to find the y intercept to put it into slope intercept form. For the first one I have the point (1,2) and the slope is undefined. For the second one I have the point (7,5) and the slope is 0. Help please!

OpenStudy (mertsj):

If the slope is undefined, it is a vertical line.

OpenStudy (mertsj):

If the slope is 0, it is a horizontal line.

OpenStudy (anonymous):

Ok so what would the first one be for the y intercept since it never touches it? or would the equation just be y=x?

OpenStudy (mertsj):

Can you graph the point (1,2) ??

OpenStudy (anonymous):

Yes it puts it on over 1 up two. its a vertical line so no y intercept and slope is undefined so I don't know what to put there

OpenStudy (mertsj):

0y+x=1 0y=-x+1

OpenStudy (mertsj):

No y intercept

OpenStudy (anonymous):

so the answer is both of them?

OpenStudy (mertsj):

1. Question: What is the y intercept? Answer: There is no y intercept. Question: What is the equation of the line in slope-intercept form? Answer: 0y=-x+1 2. Question: What is the y intercept Answer: The y intercept is 6 Question: What is the equation of the line in slope-intercept form? Answer: y = 0x+5